OpenStudy (anonymous):

solve (0,2pi] sin x sin 2x + sin 2x =0

OpenStudy (campbell_st):

well factor out sin(2x) so you have sin(2x)(sin(x) + 1) = 0 so find where sin(2x) = 0 and also find where sin(x) + 1 = 0 for the given domain.

OpenStudy (anonymous):

more explain plzz

OpenStudy (campbell_st):

ok... so just to check the question is \[\sin(x)\sin(2x) + \sin(2x) = 0\] you can factor out sin(2x) and doing so you get sin(2x)(sin(x) + 1) = 0 so the factors are sin(2x) and (sin(x) + 1) if either factor is zero, then the equation is zero. so find the values of x that make sin(2x) = 0 and sin(x) + 1 = 0 does that make sense...?

OpenStudy (anonymous):

yes it is sinx=0 sin x = -1 right

OpenStudy (campbell_st):

no sin(2x) = 0... so it means that the graph repeats 2twice over the domain of 2pi... so where is sin(2x) = 0  and yes sin(x) = -1 what angle gives that value